OpenStudy (anonymous):

Multiply and Simplify:

OpenStudy (anonymous):

\[-2\sqrt{3x}(2\sqrt{13x}-2\sqrt{11})\]

OpenStudy (anonymous):

\[-4\sqrt{39}x + 4\sqrt{33x}\] \[4\sqrt{3}(\sqrt{11x} - \sqrt{13}x)\]

OpenStudy (anonymous):

note that \[\sqrt{13x}\sqrt{3x} = \sqrt{13}\sqrt{3}\sqrt{x}\sqrt{x} =\sqrt{39}x where \sqrt{x}\sqrt{x}=x\]